# When Deep Learning Meets Polyhedral Theory: A Survey

### v3

@@ 3.5.1 Analytical results
 Telgarsky (2015) $2^{\frac{L-3}{2}}$, where $n_{i}=1$ for $i$ odd, $n_{i}=2$ for $i$ even, and $L-3$ divides by 2
@@
In the case of deep networks, where $L>1$, we need to consider how the linear regions defined up to a given layer of the network can be further partitioned by the next layers. We start by assuming that every linear region defined by the first $l-1$ layers is then subdivided into the maximum possible number of linear regions defined by the activation hyperplanes of layer $l$. That leads to the bound in Raghu et al. (2017), which is implicit in their proof of an asymptotic bound of $O(n^{n_{0}L})$, where $n$ is used as the width of every layer. However, there are many ways in which this bound can be refined upon careful examination. First, the dimension of the input of layer $l$ —i.e., the output of layer $l-1$— within each linear region is never larger than the smallest dimension among layers $1$ to $l$, since for every linear region we have an affine transformation between inputs and outputs of each layer (Montúfar, 2017). Second, the dimension of the input coming through each linear region is in fact bounded by the smallest number of active units in each of the previous layers (Serra et al., 2018). This leads to a tight upper bound for $n_{0}=1$, since it matches the lower bound in  Serra et al. (2018). Finally, the activation hyperplane of some units may not partition the linear regions because all possible inputs to the unit are in the same half-space, and in some of those cases the unit may never produce a positive output. For the number $k$ of active units in a given layer $l$, we can use the network parameters to calculate the maximum number of units that can be active in the next layer, $\mathcal{A}_{l}(k)$, as well as the number of units that can be active or inactive for different inputs, $\mathcal{I}_{l}(k)$ (Serra and Ramalingam, 2020).
@@ 3.5.2 Counting linear regions
Counting the actual number of linear regions of a given network has been a more challenging topic to explore. Serra et al. (2018) have shown that the linear regions of a trained network can be enumerated as the solutions of an MILP formulation, which has been slightly corrected in Cai et al. (2023) [^1] The MILP formulation of neural networks is discussed in Section 4.. However, MILP solutions are generally counted one by one (Danna et al., 2007), with exception of special cases (Serra and Hooker, 2020) and small subproblems (Serra, 2020), which makes this approach impractical for large neural networks. Serra and Ramalingam (2020) have shown that approximate model counting methods, which are commonly used to count the number of feasible assignments in propositional satisfiability, can be easily adapted to solution counting in MILP, which leads to an order-of-magnitude speedup in comparison with exact counting. This type of approach is particularly suitable for obtaining probabilistic lower bounds, which can complement the analytical upper bounds for the maximum number of linear regions. In Craighero et al. (2020a) and Craighero et al. (2020b), a directed acyclic graph is used to model the sets of active neurons on each layer and how they connected with those in subsequent layers. Yang et al. (2020) describe a method for decomposing the input space of rectifier networks into their linear regions by representing each linear region in terms of its face lattice, upon which the splitting operations corresponding to the transformations performed by each layer can be implemented. As the number of linear regions grows, these splitting operations can be processed in parallel. Yang et al. (2021) extend that method to convolutional neural networks. Moreover, Wang (2022) describes an algorithm for enumerating linear regions that counts adjacent linear regions with same corresponding affine function as a single linear region.
@@ 3.6.4 Function approximation
This topic also relates to the broad literature on neural networks as universal function approximators, to which the concept of linear regions helps articulating ideal conditions. As observed by Mhaskar and Poggio (2020), the optimal number of linear regions in a neural network —or, correspondingly, of pieces of the piecewise linear function modeled by it— depends on the function being approximated. In addition, linear regions were also used explicitly to build function approximations. Kumar et al. (2019) have shown that rectifier networks can be approximated to arbitrary precision with two hidden layers, the largest of which having a neuron corresponding to each different activation pattern of the original network; an exact counterpart of this result was later presented by Villani and Schoots (2023). Fan et al. (2023a) described the transformation between sufficiently wide and deep networks while arguing that the fundamental measure of complexity should be counting simplices within linear regions. In subsequent work, Fan et al. (2023b) empirically observed that linear regions tend to have a small number of higher-dimensional faces, or facets. More recently, Safran et al. (2024) studied the number of neurons necessary to approximate the maximum function.
@@
In most of the discussion so far, we have seen the benefits of using neural networks for compactly representing—and computing—piecewise linear functions and piecewise linear approximations of arbitrary functions. We would like to acknowledge an interesting line for future work posited by one of the anonymous reviewers of this survey, and which would go in the opposite direction: what functions cannot be efficiently represented by a neural network, even if they can be efficiently computed by other means?
@@ 4.1.1 Neural network verification
Refer to caption+Refer to caption=Refer to caption×$\times(\epsilon=0.15)$
@@
Consider a classification network $f:[0,1]^{n_{0}}\to\mathbb{R}^{d}$ where the $j$-th output, $f(x)_{j}$, corresponds to the probability that input $x$ is of class $j$ [^2] In actuality, we will instead typically work with the outputs corresponding to “logits”, or unnormalized probabilities. These are typically fed into a softmax layer that normalizes them to correspond to a probability distribution over the classes. This nonlinear softmax transformation is not piecewise linear, but, thankfully, it can be omitted in the context of the verification task w.l.o.g. Then consider a labeled image $\hat{x}$ known to be of class $i$, and a “target” adversarial class $k\neq i$. Then verifying local robustness of the prediction corresponds to checking $x\in\{x:||x-\hat{x}||\leq\epsilon\}\Rightarrow y=f(x)\in\{y:y_{i}\geq y_{k}\}$, where $\epsilon>0$ is a constant which prescribes the radius around which $\hat{x}$ we will search for an adversarial example.
@@
s.t. $||x-\hat{x}||\leq\epsilon.$
Any feasible solution $x$ to this problem with positive cost is an adversarial example: it is very “close” to $\hat{x}$ which has true label $i$, yet the network believes it is more likely to be of class $k$ [^3] Alternative objectives may allow us to strengthen this statement to say that the network will classify $x$ to be of class $k$. However, this will require a more complex reformulation to model this problem via MILP, so we omit it for simplicity. If, on the other hand, it is proven that the optimal objective value is negative, this proves that $f$ is robust (at least in the neighborhood around $\hat{x}$). Note that the verification problem can terminate once the sign of the optimal objective value is determined, but solving the problem returns an optimal adversarial example.
@@ 4.1.2 Neural network as proxy
Another situation in which one may want to solve an optimization problem containing trained neural networks is when one would like to optimize some other, unknown function for which historical input/output data exist. A similar situation arises when solving an optimization problem where (some of) the constraints are overly complicated, but samples from the constraints can be queried and used to train a simpler surrogate model. In these cases, a neural network can be trained to approximate this underlying, unknown or complicated function. Then, since the neural network is known, the resulting optimization problem is a deterministic piecewise linear problem. Note that we focus here on using a neural network as a surrogate; neural networks can additionally learn other components of an optimization problem, e.g., uncertainty sets for robust optimization (Goerigk and Kurtz, 2023).
Several software tools have been developed for this class of problems. For the case of constraint learning, JANOS (Bergman et al., 2022) and OptiCL (Maragno et al., 2023) both provide functionality for learning a ReLU neural network to approximate a constraint based on data and embedding the learned neural network in MILP. The reluMIP package (Lueg et al., 2021) has also been introduced to handle the latter embedding step. More generally, OMLT (Ceccon et al., 2022) translates neural networks to pyomo optimization blocks, including various MILP formulations and activation functions. Finally, recent developments in gurobipy [^4] https://github.com/Gurobi/gurobi-machinelearning enable directly parsing in trained neural networks.
@@
Chemical engineering also presents applications where surrogate models have proven beneficial for optimization, as is the subject of recent reviews (Bhosekar and Ierapetritou, 2018, McBride and Sundmacher, 2019, Tsay and Baldea, 2019). In particular, ReLU neural networks can be seamlessly embedded in larger MILP problems such as flow networks and reservoir control where the other constraints are also mixed-integer linear (Grimstad and Andersson, 2019, Say et al., 2017, Yang et al., 2022). Focusing on control applications where the neural network is embedded in a MILP that must be solved repeatedly, Katz et al. (2020) use multiparametric programming to learn the solution map of the resulting MILP itself, which is also piecewise affine. An emerging area of research uses verification tools to reason about neural networks used as controllers, e.g., see Johnson et al. (2020). These applications involve optimization formulations combining the neural network with constraints defining the controlled system. For example, verification can be used to bound the reachable set (Sidrane et al., 2022, Sosnin and Tsay, 2024) (alongside piecewise linear bounds on the dynamical system) or the maximum error against a baseline controller (Schwan et al., 2023).
Finally, applications for optimization over neural networks arise in machine learning applications. MILP formulations can be used to compress neural networks (Serra et al., 2020, 2021, ElAraby et al., 2020), which consequently result in more tractable surrogate models (Kody et al., 2022). The main idea is to use MILP to identify stable nodes, i.e., nodes that are always on or off over an input domain, which can then be algebraically eliminated. Optimization has also been employed in techniques for feature selection, based on identifying strongest input nodes (Sildir and Aydin, 2022, Zhao et al., 2023). In the context of Bayesian optimization, Volpp et al. (2020) use reinforcement learning to meta-learn acquisition functions parameterized as neural networks; selecting ensuing query points then requires optimization over the trained acquisition function. Later work modeled both the acquisition function and feasible region in black-box optimization as neural networks (Papalexopoulos et al., 2022). In that work, exploration and exploitation are balanced via Thompson sampling and training multiple neural networks from a random parameter initialization.
@@ 4.2.1 Single neuron relaxations
gr =$=\Set{({\bm{h}}^{l-1},h^{l}_{i})\in\mathcal{D}^{l-1}\times\mathbb{R}}{h^{l}_{i}=0\geq{\bm{w}}^{l}_{i}{\bm{h}}^{l-1}+b^{l}_{i}}$
$\cup\Set{({\bm{h}}^{l-1},h^{l}_{i})\in\mathcal{D}^{l-1}\times\mathbb{R}}{h^{l}_{i}={\bm{w}}^{l}_{i}{\bm{h}}^{l-1}+b^{l}_{i}\geq 0}.$
@@
This also generalizes in a straightforward manner to more complex feedforward network architectures (e.g. convolutions, or sparse or skip connections), though we omit the explicit description for notational simplicity.
@@ 4.2.2 Beyond the scope of this survey
The effectiveness of the single-neuron formulations described above is bounded by the tightness of the optimal univariate formulation; this property is known as the “single-neuron barrier” (Salman et al., 2019). This has motivated research in convex relaxations that jointly account for multiple neurons within a layer (Singh et al., 2019a). Nevertheless, the analysis of polyhedral formulations for multiple neurons simultaneously quickly becomes intractable, and is beyond the scope of this survey. Instead, we point the interested reader to the recent survey by Roth (2021), and highlight a few approaches taken in the literature. Multi-neuron analysis has been used to: improve bounds tightening schemes (Rössig and Petkovic, 2021), prune linearizable neurons (Botoeva et al., 2020), design dual decompositions (Ferrari et al., 2022), and generate strengthening inequalities (Serra and Ramalingam, 2020). Similarly, we do not review formulations for ensembles of ReLU networks, though MILP formulations have been proposed (Wang et al., 2021, 2023).
Additionally, recent works have exploited polyhedral structure to develop sampling-based strategies, which can be used to warm-start MILP (i.e., primal heuristics) or accelerate local search in verification (Perakis and Tsiourvas, 2022, Wu et al., 2022). Lombardi et al. (2017) computationally compare MILP against local search and constraint programming approaches. In a related vein, Cheon (2022) examines local solutions and proposes an outer approximation method to improve gradient-based optimization. Finally, following Raghunathan et al. (2018), a large body of work has presented optimization-based methods for verification that use semidefinite programming concepts (Dathathri et al., 2020, Fazlyab et al., 2020, Newton and Papachristodoulou, 2021). Notably, Batten et al. (2021) showed how combining semidefinite and MILP formulations can produce a new formulation that is tighter than both. This was later extended with reformulation-linearization technique, or RLT, cuts (Lan et al., 2022). While related to linear programming and other methods based on convex relaxations, this stream of work is beyond the scope of this survey. We refer the reader to Zhang (2020) for a discussion of the tightness of these formulations.
@@ 4.3.1 The big-$M$ formulation
It is worth dwelling on where this lack of strength comes from. If the input ${\bm{h}}^{l-1}$ is one dimensional, the big-$M$ formulation is locally ideal (Vielma, 2015): the extreme points of the linear programming relaxation (4a-4d) naturally satisfy the integrality constraints (4e). However, this fails to hold in the general multivariate input case. To see why, observe that the bounds on the input variables ${\bm{h}}^{l-1}$ are only coupled with the logic involving the binary variable $z$ only in an aggregated sense, through the coefficients $M^{l}_{i,-}$ and $M^{l}_{i,+}$. In other words, the “shape” of the pre-activation domain is not incorporated directly into the big-$M$ formulation. Furthermore, the strength of this formulation highly depends on the big-$M$ coefficients. These coefficients can be obtained using techniques ranging from basic interval arithmetic to optimization-based bounds tightening. Grimstad and Andersson (2019) show how constraints external to the neural network can yield tighter bounds via optimization- or feasibility-based bounds tightening. Rössig and Petkovic (2021) compare several methods for deriving bounds and further develop optimization-based bounds tightening based on pairwise dependencies between variables.

### v4

@@ 3.5.1 Analytical results
 Telgarsky (2015) $2^{\frac{L-3}{2}}$, where $n_{i}=1$ for $i$ odd, $n_{i}=2$ for $i$ even, and L3L-3 divides by 2
@@
In the case of deep networks, where $L>1$, we need to consider how the linear regions defined up to a given layer of the network can be further partitioned by the next layers. We start by assuming that every linear region defined by the first l1l-1 layers is then subdivided into the maximum possible number of linear regions defined by the activation hyperplanes of layer $l$. That leads to the bound in Raghu et al. (2017), which is implicit in their proof of an asymptotic bound of $O(n^{n_{0}L})$, where $n$ is used as the width of every layer. However, there are many ways in which this bound can be refined upon careful examination. First, the dimension of the input of layer $l$ —i.e., the output of layer l1l-1— within each linear region is never larger than the smallest dimension among layers $1$ to $l$, since for every linear region we have an affine transformation between inputs and outputs of each layer (Montúfar, 2017). Second, the dimension of the input coming through each linear region is in fact bounded by the smallest number of active units in each of the previous layers (Serra et al., 2018). This leads to a tight upper bound for $n_{0}=1$, since it matches the lower bound in  Serra et al. (2018). Finally, the activation hyperplane of some units may not partition the linear regions because all possible inputs to the unit are in the same half-space, and in some of those cases the unit may never produce a positive output. For the number $k$ of active units in a given layer $l$, we can use the network parameters to calculate the maximum number of units that can be active in the next layer, $\mathcal{A}_{l}(k)$, as well as the number of units that can be active or inactive for different inputs, $\mathcal{I}_{l}(k)$ (Serra and Ramalingam, 2020).
@@ 3.5.2 Counting linear regions
Counting the actual number of linear regions of a given network has been a more challenging topic to explore. Serra et al. (2018) have shown that the linear regions of a trained network can be enumerated as the solutions of an MILP formulation, which has been slightly corrected in Cai et al. (2023) [^1] The MILP formulation of neural networks is discussed in Section 4.. However, MILP solutions are generally counted one by one (Danna et al., 2007), with exception of special cases (Serra and Hooker, 2020) and small subproblems (Serra, 2020), which makes this approach impractical for large neural networks. In fact, Stargalla et al. (2025) have shown that counting linear regions, as we define them in this survey, is #P-complete. Serra and Ramalingam (2020) have shown that approximate model counting methods, which are commonly used to count the number of feasible assignments in propositional satisfiability, can be easily adapted to solution counting in MILP, which leads to an order-of-magnitude speedup in comparison with exact counting. This type of approach is particularly suitable for obtaining probabilistic lower bounds, which can complement the analytical upper bounds for the maximum number of linear regions. In Craighero et al. (2020a) and Craighero et al. (2020b), a directed acyclic graph is used to model the sets of active neurons on each layer and how they connected with those in subsequent layers. Yang et al. (2020) describe a method for decomposing the input space of rectifier networks into their linear regions by representing each linear region in terms of its face lattice, upon which the splitting operations corresponding to the transformations performed by each layer can be implemented. As the number of linear regions grows, these splitting operations can be processed in parallel. Yang et al. (2021) extend that method to convolutional neural networks. Moreover, Wang (2022) describes an algorithm for enumerating linear regions that counts adjacent linear regions with same corresponding affine function as a single linear region.
@@ 3.6.4 Function approximation
This topic also relates to the broad literature on neural networks as universal function approximators, to which the concept of linear regions helps articulating ideal conditions. As observed by Mhaskar and Poggio (2020), the optimal number of linear regions in a neural network —or, correspondingly, of pieces of the piecewise linear function modeled by it— depends on the function being approximated. In addition, linear regions were also used explicitly to build function approximations. Kumar et al. (2019) have shown that rectifier networks can be approximated to arbitrary precision with two hidden layers, the largest of which having a neuron corresponding to each different activation pattern of the original network; an exact counterpart of this result based on taking the limit of arbitrarily large parameters was later presented by Villani and Schoots (2023). A more recent development by Zanotti (2025) have shown that two hidden layers suffice to represent a piecewise linear function with a number of neurons linear on the number of linear regions. Fan et al. (2023a) described the transformation between sufficiently wide and deep networks while arguing that the fundamental measure of complexity should be counting simplices within linear regions. In subsequent work, Fan et al. (2023b) empirically observed that linear regions tend to have a small number of higher-dimensional faces, or facets. More recently, Safran et al. (2024) studied the number of neurons necessary to approximate the maximum function.
@@
In most of the discussion so far, we have seen the benefits of using neural networks for compactly representing—and computing—piecewise linear functions and piecewise linear approximations of arbitrary functions. We would like to acknowledge an interesting line for further work posited by one of the anonymous reviewers of this survey, and which would go in the opposite direction: what functions cannot be efficiently represented by a neural network, even if they can be efficiently computed by other means? For example, Hertrich and Loho (2024) establish lower bounds on the size of certain types of neural networks that solve linear programming problems.
@@ 4.1.1 Neural network verification
[…]
@@
[…]
@@
[…]
Any feasible solution $x$ to this problem with positive cost is an adversarial example: it is very “close” to $\hat{x}$ which has true label $i$, yet the network believes it is more likely to be of class kk. [^3] Alternative objectives may allow us to strengthen this statement to say that the network will classify $x$ to be of class $k$. However, this will require a more complex reformulation to model this problem via MILP, so we omit it for simplicity. If, on the other hand, it is proven that the optimal objective value is negative, this proves that $f$ is robust (at least in the neighborhood around $\hat{x}$). Note that the verification problem can terminate once the sign of the optimal objective value is determined, but solving the problem returns an optimal adversarial example.
@@ 4.1.2 Neural network as proxy
Another situation in which one may want to solve an optimization problem containing trained neural networks is when one would like to optimize some other, unknown function for which historical input/output data exist. A similar situation arises when solving an optimization problem where (some of) the constraints are overly complicated, but samples from the constraints can be queried and used to train a simpler surrogate model. In these cases, a neural network can be trained to approximate this underlying, unknown or complicated function. Then, since the neural network is known, the resulting optimization problem is a deterministic piecewise linear problem. Note that we focus here on using a neural network as a surrogate; neural networks can additionally learn other components of an optimization problem, e.g., uncertainty sets for robust optimization (Goerigk and Kurtz, 2023) or second-stage value functions for two-stage stochastic programming (Alcántara et al., 2025, Patel et al., 2022).
Several software tools have been developed for this class of problems. For the case of constraint learning, JANOS (Bergman et al., 2022) and OptiCL (Maragno et al., 2023) both provide functionality for learning a ReLU neural network to approximate a constraint based on data and embedding the learned neural network in MILP. The reluMIP package (Lueg et al., 2021) has also been introduced to handle the latter embedding step. More generally, OMLT (Ceccon et al., 2022) translates neural networks to pyomo optimization blocks, including various architectures (e.g,. CNNs, GNNs), MILP formulations, and activation functions. More recently, PySCIPOpt-ML (Turner et al., 2025) and MathOptAI.jl (Dowson et al., 2025) provide similar interfaces for SCIP and JuMP, respectively. Finally, recent developments in gurobipy [^4] https://github.com/Gurobi/gurobi-machinelearning enable directly parsing in trained neural networks.
@@
Chemical engineering also presents applications where surrogate models have proven beneficial for optimization, as is the subject of recent reviews (Bhosekar and Ierapetritou, 2018, McBride and Sundmacher, 2019, Tsay and Baldea, 2019). In particular, ReLU neural networks can be seamlessly embedded in larger MILP problems such as flow networks and reservoir control where the other constraints are also mixed-integer linear (Grimstad and Andersson, 2019, Say et al., 2017, Yang et al., 2022, Liu and Dvorkin, 2025). Other recent uses for ReLU neural networks within MILP models are for representing the uncertainty in response variables in stochastic optimization (Alcántara and Ruiz, 2023) and as a proxy for solving routing problems associated with a facility location problem (Kaleem and Subramanyam, 2024). Focusing on control applications where the neural network is embedded in a MILP that must be solved repeatedly, Katz et al. (2020) use multiparametric programming to learn the solution map of the resulting MILP itself, which is also piecewise affine. An emerging area of research uses verification tools to reason about neural networks used as controllers, e.g., see Johnson et al. (2020). These applications involve optimization formulations combining the neural network with constraints defining the controlled system. For example, verification can be used to bound the reachable set (Sidrane et al., 2022, Sosnin and Tsay, 2024) (alongside piecewise linear bounds on the dynamical system) or the maximum error against a baseline controller (Schwan et al., 2023).
Finally, applications for optimization over neural networks arise in machine learning applications. MILP formulations can be used to compress neural networks (Serra et al., 2020, 2021, ElAraby et al., 2020), which consequently result in more tractable surrogate models (Kody et al., 2022). The main idea is to use MILP to identify stable nodes, i.e., nodes that are always on or off over an input domain, which can then be algebraically eliminated. Optimization has also been employed in techniques for feature selection, based on identifying strongest input nodes (Sildir and Aydin, 2022, Zhao et al., 2023). In the context of Bayesian optimization, Volpp et al. (2020) use reinforcement learning to meta-learn acquisition functions parameterized as neural networks; selecting ensuing query points then requires optimization over the trained acquisition function. Later work modeled both the acquisition function and feasible region in black-box optimization as neural networks (Papalexopoulos et al., 2022). In that work, exploration and exploitation are balanced via Thompson sampling and training multiple neural networks from a random parameter initialization. Another application is the generation of counterfactual explanations for the neural network (Kanamori et al., 2021).
@@ 4.2.1 Single neuron relaxations
[…]
@@
This also generalizes in a straightforward manner to more complex feedforward network architectures (e.g. convolutions, or sparse or skip connections), though we omit the explicit description for notational simplicity. One interesting line of work here explores formulations specific to graph neural networks (GNNs), using MILP techniques to simultaneously handle the GNN and the combinatorial input search space over graphs (McDonald et al., 2024, Zhang et al., 2023b).
@@ 4.2.2 Beyond the scope of this survey
The effectiveness of the single-neuron formulations described above is bounded by the tightness of the optimal univariate formulation; this property is known as the “single-neuron barrier” (Salman et al., 2019). This has motivated research in convex relaxations that jointly account for multiple neurons within a layer (Singh et al., 2019a). Nevertheless, the analysis of polyhedral formulations for multiple neurons simultaneously quickly becomes intractable, and is beyond the scope of this survey. Instead, we point the interested reader to the recent survey by Roth (2021), and highlight a few approaches taken in the literature. Multi-neuron analysis has been used to: improve bounds tightening schemes (Rössig and Petkovic, 2021), prune linearizable neurons (Botoeva et al., 2020), design dual decompositions (Ferrari et al., 2022), and generate strengthening inequalities (Serra and Ramalingam, 2020, Zhou et al., 2024). Similarly, we do not review formulations for ensembles of ReLU networks, though MILP formulations have been proposed (Wang et al., 2021, 2023).
Additionally, recent works have exploited polyhedral structure to develop sampling-based strategies, which can be used to warm-start MILP (i.e., primal heuristics) or accelerate local search in verification (Perakis and Tsiourvas, 2022, Tong et al., 2024, Wu et al., 2022). Other recent approaches include the use of surrogate models based on pruned neural networks (Pham et al., 2025) and dualizing the constraints associated with ReLU in the optimization model (Liu and Dvorkin, 2025). Lombardi et al. (2017) computationally compare MILP against local search and constraint programming approaches. Plate et al. (2025) analyze the impact of a variety of factors on the runtimes of MILP with embedded neural networks. In a related vein, Cheon (2022) examines local solutions and proposes an outer approximation method to improve gradient-based optimization. Finally, following Raghunathan et al. (2018), a large body of work has presented optimization-based methods for verification that use semidefinite programming concepts (Dathathri et al., 2020, Fazlyab et al., 2020, Newton and Papachristodoulou, 2021). Notably, Batten et al. (2021) showed how combining semidefinite and MILP formulations can produce a new formulation that is tighter than both. This was later extended with reformulation-linearization technique, or RLT, cuts (Lan et al., 2022). While related to linear programming and other methods based on convex relaxations, this stream of work is beyond the scope of this survey. We refer the reader to Zhang (2020) for a discussion of the tightness of these formulations. More recently, Karia et al. (2025) studied the embedding of Kolmogorov-Arnold networks (Liu et al., 2025), which may have activation functions requiring a nonlinear optimization formulation.
@@ 4.3.1 The big-$M$ formulation
It is worth dwelling on where this lack of strength comes from. If the input ${\bm{h}}^{l-1}$ is one dimensional, the big-$M$ formulation is locally ideal (Vielma, 2015): the extreme points of the linear programming relaxation (4a-4d) naturally satisfy the integrality constraints (4e). However, this fails to hold in the general multivariate input case. To see why, observe that the bounds on the input variables ${\bm{h}}^{l-1}$ are only coupled with the logic involving the binary variable $z$ only in an aggregated sense, through the coefficients $M^{l}_{i,-}$ and $M^{l}_{i,+}$. In other words, the “shape” of the pre-activation domain is not incorporated directly into the big-$M$ formulation. Furthermore, the strength of this formulation highly depends on the big-$M$ coefficients. These coefficients can be obtained using techniques ranging from basic interval arithmetic to optimization-based bounds tightening. Grimstad and Andersson (2019) show how constraints external to the neural network can yield tighter bounds via optimization- or feasibility-based bounds tightening. Rössig and Petkovic (2021) compare several methods for deriving bounds and further develop optimization-based bounds tightening based on pairwise dependencies between variables. Zhou et al. (2024) present a rolling-horizon decomposition for bounds tightening based on solving a subproblem for each layer of the network.